# Ab initio recombination in the expanding ultracold plasmas

### v1

@@ II.4 Initial conditions
An accurate specification of the initial conditions for ultracold plasmas is a rather nontrivial tack. For example, if the plasma is produced by instantaneous photoionization of neutral atoms, the initial electron positions should be substantially correlated with the ionic ones, and the distribution of their velocities will be non-Gaussian Niffenegger, Gilmore, and Robicheaux (2011). In the opposite case, when the ionization process takes some time, the released electrons will be mixed in space between the ions and approximately thermalized, resulting in the uniform coordinate and Gaussian velocity distributions. Since we are interested in the recombination events occurring after a sufficiently long temporal evolution, exact specification of the initial conditions should not be important. So, for the sake of simplicity, we shall follow the second option. Namely, initial positions of the electrons $r^{*}_{i0}$ and ions $R^{*}_{i0}$ ($i=1,\dots,N$) will be given by random numbers distributed uniformly within the box, $L^{*}/2\leq(x^{*},y^{*},z^{*})\leq L^{*}/2$.
@@
On the other hand, the initial electron velocities ${\dot{r}}^{*}_{i0}$ will be specified by the normal (Gaussian) distribution with a root-mean-square deviation ${\sigma}_{v0}^{*}$ over each Cartesian coordinate. This quantity evidently characterizes the initial electron temperature. As was shown in our work Dumin (2011), the formal definition $T_{e}=(2/3)\,K_{e}/k_{\rm B}$ (where $K_{e}$ is the electron kinetic energy, and $k_{\rm B}$ is the Boltzmann constant) should work rather well even for the electrons strongly interacting with ions. At last, as was already mentioned above, the ions are assumed to move by inertia due to their very large masses, i.e., in the co-moving reference frame, $R^{*}_{i}=\mbox{const}$ and ${\dot{R}}^{*}_{i}=0$.
@@ III Results of the Simulations
The simulations were performed for various numbers of the charged particles of each kind $N$ in the basic cell, from 10 to 1000. Since the employed algorithm of numerical integration—without the adaptive stepsize control—was computationally expensive, a sufficiently long interval of integration (about $1000\,\tau$) was achieved only at the small number of particles (e.g., $N=10$); and the most convincing patterns of recombination were obtained just in this case. On the other hand, when $N$ was substantially larger, the achievable interval of integration was a few times shorter; so that only the initial stage of recombination was observed. (To avoid misunderstanding, let us remind once again that the actual number of particles taken into account in calculation of the Coulomb forces, including the “mirror” cells, was much greater: for example, even at $N=10$ it varied from a few hundred thousand to a few million, depending on the rate of convergence of the Coulomb sums at each particular step of integration.) In the results presented below, we used the initial dispersion of the electron velocities ${\sigma}_{v0}^{*}=3.0$. In other words, the plasma was originally weakly non-ideal, since its initial kinetic energy exceeded the potential one by an order of magnitude. The overall plasma expansion rate was taken to be $u_{0}^{*}=0.1$.
@@
Figure 3: Temporal dependence of the kinetic (red curves) and potential (blue curves) energies of all particles in the basic cell at various “magnifications” (successive panels, from top to bottom). The regions of subsequent magnification are marked by the rectangles.
Figure 3 shows a temporal behavior of the total kinetic and potential energies of all particles in the basic cell, including the energy of interaction with the “mirror” images. As can be expected, in the upper panel one can see initially a sharp decay in the kinetic energy, which is just an adiabatic cooling of the expanding plasma. A much more interesting and nontrivial feature is that this cooling is associated with a development of fast oscillations (a series of the sharp equidistant peaks), surviving up to the end of simulation (which was about two times longer than the time interval presented in figure). Moreover, starting from $250{-}300~\tau$ one can see onset of the second, much larger period of modulation, which also survives up to the end of our simulation. The presence of these two periods is clearly seen in the second, enlarged panel. At last, inspection of the third, most magnified panel gives us a hint that the modulation of the kinetic and potential energies is caused just by the passages of an electron, captured into a strongly elliptical orbit, near the pericenter and apocenter of its trajectory.
@@
Figure 4: Trajectory of the first (short-period) captured electron viewed in three coordinate planes at two different magnifications (three top and three bottom panels, respectively). The region of formation of a captured state of the electron is marked by the red dotted circle. The coordinate scales in various directions are different.
@@
Figure 5: Trajectory of the second (long-period) captured electron viewed in three coordinate planes. The region of formation of a captured state of the electron is marked by the red dotted circle. The coordinate scales in various directions are different.
@@
A more careful analysis of the electron trajectories, presented in Figs. 4 and 5, confirms this conjecture. It is seen that the above-mentioned electrons—after some periods of a “random walk”—really form the confined states in the co-moving reference frame. So, they should result in the recombination with the corresponding ions if quantum effects will be taken into account.
The trajectories of other eight electrons in the basic cell do not exhibit any features of localization. Consequently, we can estimate the total efficiency of recombination as 20%. This is actually the main result of our simulations.
@@ IV Discussion
Although the employed sample of particles was very small (10 electrons and 10 ions in the basic cell), the results obtained are in agreement with the experimental measurements Killian et al. (2001) as well as with the earlier modeling Niffenegger, Gilmore, and Robicheaux (2011). In the last-cited paper, the criterion of recombination was introduced rather formally as the percentage of electrons located closer than the specified distance (e.g., 0.2 Wigner–Seitz radius) to the nearest ion. Surprisingly, this semi-empirical criterion gave approximately the same result, $17{-}18\,\%$, as our ab initio approach.
@@
Yet another widely-used criterion Lankin and Norman (2009a) is to find the electrons that performed a specified minimal number of revolutions (e.g., 1, 2, 4, or 6) about the nearest ion and to assume that they will ultimately recombine to the neutral atoms. In fact, this approach is even more popular and often considered as justified better. Unfortunately, analysis of the results of our simulations did not provide any significant support in favor of this criterion. So, this subject still needs to be studied in more detail.
As was already mentioned above, the number of particles in the basic cell simulated in detail by our method, $N=10$, was rather small. This was because each computational run on the ordinary PC, without a parallelization, took a few months. So, when the number of particles was much greater, e.g. 100 or 1000, we were able to perform the simulation only at the considerably shorter time intervals, about $100\,\tau$, when the recombination process was not yet complete.
The low speed of calculations was caused by the fact that we employed neither the adaptive stepsize control for integration, nor Ewald summation technique for the Coulomb sums. Unfortunately, when these methods were used in our earlier attempts, we were unable to reproduce the captured electron trajectories at all, most probably, because of the additional errors introduced by the corresponding procedures. So, we used here a more straightforward approach. However, one can expect that it will be possible in future to introduce the adaptive stepsize control and Ewald summation into our algorithm of the scalable (co-moving) reference frame, which should make this approach much more efficient computationally.
@@ V Conclusions
    To the best of our knowledge, the performed work is the first successful attempt of ab initio simulation of recombination in the expanding ultracold plasmas, which clearly demonstrates how some of electrons become localized and form the stable compact pairs with the respective ions.
@@
    A cornerstone of our approach is employing the scalable reference frame, co-moving with the expanding plasma. Such transformation formally converts the effect of overall plasma expansion into the effective frictional force in the fixed coordinate frame. This enables us both to avoid some computational problems and to simplify interpretation of the results of simulation.
@@
    The efficiency of recombination obtained by this method (about 20% for the realistic experimental conditions) turns out to be in agreement both with the values derived by some semi-empirical approaches Niffenegger, Gilmore, and Robicheaux (2011) and with the laboratory measurements Killian et al. (2001).
@@
    Unfortunately, the proposed modeling scheme is rather expensive computationally. Therefore, it is primarily intended for testing the accuracy of other (semi-empirical) methods rather than for a routine usage. However, it is expected that the efficiency of this approach can be increased in future by introducing new versions of the adaptive stepsize control for the numerical integration and the refined Ewald summation procedures for calculation of the Coulomb sums.
@@ Acknowledgements.
YVD is grateful to A.A. Bobrov, P.R. Levashov, S.A. Mayorov, J.-M. Rost, S.A. Saakyan, U. Saalmann, V.S. Vorob’ev, B.B. Zelener, and B.V. Zelener for fruitful discussions and valuable suggestions.

### v2

@@ II.5 Initial conditions
An accurate specification of the initial conditions for ultracold plasmas is a rather nontrivial tack. For example, if the plasma is produced by instantaneous photoionization of neutral atoms, then the initial electron positions should be substantially correlated with the ionic ones, and distribution of the electron velocities will be non-Gaussian Niffenegger, Gilmore, and Robicheaux (2011). In the opposite case, when the ionization process takes some time, the released electrons will be mixed in space between the ions and approximately thermalized, resulting in the uniform coordinate and Gaussian velocity distributions. Since we are interested in the recombination events occurring after a quite long temporal evolution, exact specification of the initial conditions should not be important. So, for the sake of simplicity, we followed the second option. Namely, initial positions of the electrons $r^{*}_{i0}$ and ions $R^{*}_{i0}$ ($i=1,\dots,N$) were given by the random numbers distributed uniformly within the box, $L^{*}/2\leq(x^{*},y^{*},z^{*})\leq L^{*}/2$.
@@
On the other hand, the initial electron velocities ${\dot{r}}^{*}_{i0}$ were specified by the normal (Gaussian) distribution with a root-mean-square deviation ${\sigma}_{v0}^{*}$ over each Cartesian coordinate. This quantity evidently characterizes the initial electron temperature. As was shown in our work Dumin (2011), the formal definition $T_{e}=(2/3)\,E_{\rm kin}/k_{\rm B}$ (where $E_{\rm kin}$ is the electron kinetic energy, and $k_{\rm B}$ is the Boltzmann constant) should work rather well even if the electrons strongly interact with ions. At last, as was already mentioned above, the ions are assumed to move by inertia due to their very large masses, i.e., in the co-moving reference frame they are at rest: $R^{*}_{i}\equiv\mbox{const}$ and ${\dot{R}}^{*}_{i}\equiv 0$.
@@ III Results of the simulations
Figure 4: Temporal dependence of the kinetic (red curves) and potential (blue curves) energies of all particles in the basic cell at various “magnifications” (successive panels, from top to bottom). The regions of subsequent magnification are marked by the gray rectangles.
The simulations were performed for various numbers of the charged particles of each kind $N$ in the basic cell, from 10 to 1000. Since the employed algorithm of numerical integration without the adaptive stepsize control was computationally expensive, a sufficiently long interval of integration (about $1000\,\tau$) was achieved only for the small number of particles (e.g., $N=10$); and the most convincing patterns of recombination were obtained just in this case. On the other hand, when $N$ was substantially larger, the achievable interval became shorter by an order of magnitude; so that only the initial stage of recombination was observed, and it was difficult to derive a reliable conclusion on its total efficiency. (To avoid misunderstanding, let us remind once again that the actual number of particles taken into account in calculation of the Coulomb forces, including the “mirror” cells, was much greater: for example, even at $N=10$ it varied from a few hundred thousand to a few million, depending on the rate of convergence of the Coulomb sums at each particular step of integration.)
In the results presented below, we used the initial dispersion of the electron velocities ${\sigma}_{v0}^{*}=3.0$. In other words, the plasma was originally weakly non-ideal (${\Gamma}_{e}\sim 0.1$), since its initial kinetic energy exceeded the potential (Coulomb) energy only by one order of magnitude.
The overall plasma expansion rate was taken to be $u_{0}^{*}=0.1$. Then, “physical” size of the simulation cell (and, therefore, of the entire plasma cloud) changes in the course of time according to formula (7). Since the simulations presented below were conducted up to $t^{*}\approx 600$, the cloud expanded approximately by 60 times and its density, respectively, dropped by over 5 orders of magnitude. Such variation well corresponds to the typical experimental situations Killian et al. (1999, 2007).
@@
Figure 5: Trajectory of the first (short-period) captured electron viewed in three coordinate planes at two different magnifications (three top and three bottom panels, respectively) as well as its kinetic energy (the fourth panels in the top and bottom). The region of formation of a captured state of the electron is marked by the red dotted circles.
@@
Figure 6: Trajectory of the second (long-period) captured electron viewed in three coordinate planes (two top and left bottom panels) as well as its kinetic energy (right bottom panel). The region of formation of a captured state of the electron is marked by the red dotted circles.
Figure 4 shows a temporal behavior of the total kinetic and potential energies of all particles in the basic cell, including the energy of interaction with the “mirror” images. As can be expected, in the upper panel one can see initially a sharp decay in the kinetic energy, which is just an adiabatic cooling of the expanding plasma. As a result, the electron Coulomb coupling parameter ${\Gamma}_{e}$ quickly decays from ${\sim}0.1$ to unity and approximately stabilizes at this level.
The most interesting and nontrivial effect observed in this figure is the development of fast oscillations (a series of the sharp equidistant peaks), surviving up to the end of our simulation. Moreover, starting from $t^{*}\approx 250{-}300$ one can see onset of the second, much longer period of oscillations, which also survives up to the end of simulation. A presence of the corresponding two periods is clearly seen in the second, enlarged panel. At last, inspection of the third, most magnified panel gives us a hint that modulation of the kinetic and potential energies is caused just by the passages of an electron, captured into a strongly elliptical orbit, near the pericenter and apocenter of its trajectory.
@@
A more careful analysis of the electron trajectories, presented in Figs. 5 and 6, confirms this conjecture. It is seen that the above-mentioned electrons—after some intervals of the “random walk”—really form the confined states. Their orbits are gradually contracting just because they are depicted in the co-moving reference frame, expanding with the plasma cloud. In fact, these pairs are detached from the overall expansion and have an approximately constant size; but they look contracting in the scalable coordinate system, where size of the basic cell $L^{*}$ remains constant.
The captured electrons should ultimately recombine with the corresponding ions into the neutral atoms when quantum effects will be taken into account. This corresponds to the well-known mechanism of the collisional two-stage recombination Massey and Burhop (1956); Smirnov (1981), whose total rate is governed mostly by the first stage (i.e., formation of the classical bound pairs). The trajectories of other eight electrons in the basic cell do not exhibit any features of localization. Consequently, we can estimate the total efficiency of recombination as 20%. This is actually the main result of our simulations.
A further inspection of the trajectories shows that the first electron (Fig. 5) became captured already at $t^{*}\approx 20{-}25$ (i.e., when the plasma cloud expanded by 3–3.5 times); but its orbit initially was very irregular and, therefore, did not exhibit a definite period. Besides, it is interesting that the long-period modulation of energy, caused by the second electron, is visible not only in the energy of the entire system (Fig. 4) but also in the energy of the first electron (bottom right panel in Fig. 5). In general, this is not surprising, because the quickly-oscillating first electron should naturally experience the long-period perturbations by the second electron.
At last, inspection of the bottom right panel in Fig. 6 shows that oscillations of energy of the second electron began to develop already at $t^{*}\approx 50{-}100$, while they are visible in the total energy (top panel in Fig. 4) only starting from $t^{*}\approx 250{-}300$. A gradual decay of the amplitude of these oscillations is evidently caused by a decreasing ellipticity of the orbit.
@@ IV Discussion
In the present work we followed the well-known idea of collisional recombination Massey and Burhop (1956); Smirnov (1981) that the bound electron-ion pairs are formed initially due to the classical dynamics, and then almost all of them recombine to neutral atoms due to the quantum transitions (which are commonly assumed to occur much faster). So, the total efficiency of recombination should be equal approximately to the rate of formation of the classically-bound electron–ion pairs. (Of course, such an approach does not assume establishment of the thermodynamic equilibrium.)
The plasma in our simulations was initially weakly-coupled (${\Gamma}_{e}\sim 0.1$) but then quickly evolved to the moderately-coupled state (${\Gamma}_{e}\sim 1$), and the recombination processes developed just in this regime. Let us emphasize that the most of earlier UCP simulations were carried out in a box of fixed size for a set of “predefined” parameters (density, temperature, etc.). As distinct from that works, our method of “scalable frames” enables us to simulate the expanding plasmas by a self-consistent way, i.e., the recombination occurs at the time-dependent density, temperature, etc.; and all these parameters are determined “automatically” from the same simulation.
@@
Although the sample of particles employed in our work was rather small, the results obtained turned out to be in agreement both with the experimental measurements Killian et al. (2001) as well as with the earlier numerical modeling Niffenegger, Gilmore, and Robicheaux (2011). In fact, the criterion of recombination in the last-cited paper was introduced rather arbitrarily as the percentage of electrons located closer than 0.2 Wigner–Seitz radius to the nearest ion. However, it is interesting that such choice of the “critical” radius enabled that authors to get approximately the same total efficiency of recombination (namely, $17{-}18\,\%$) as in our ab initio simulation.
@@
Yet another frequently-used criterion Lankin and Norman (2009a) is to find the electrons that performed a specified minimal number of revolutions (e.g., 1, 2, 4, or 6) about the nearest ion and to assume that they will ultimately recombine to the neutral atoms. In fact, this approach is even more popular and often considered as justified better. Unfortunately, analysis of the results of our simulations did not provide a clear support in favor of such criterion. So, this subject still needs to be studied in more detail.
As was already mentioned above, the employed numerical algorithm is time-consuming: even at the relatively small number of particles in the basic cell, each computational run on the ordinary PC (without parallelization) took a few months. So, when the number of particles was much greater, e.g. 100 or 1000, we were able to perform simulations only for the considerably shorter time intervals, about $100\,\tau$, when the recombination processes were not complete yet. The low speed of calculations was caused by the fact that we used neither the adaptive stepsize control for integration, nor Ewald summation technique for the Coulomb sums. However, one can expect that it will be possible in future to incorporate these methods into our algorithm of the scalable (co-moving) reference frames, which should make the such simulations much faster.
In fact, it is not sufficiently clear how sensitive the recombination statistics is to the accurate treatment of the long-range forces and if the corresponding procedures can be reasonably simplified. On the one hand, the bound states cannot be evidently formed by the two-body interactions (since the total energy of the pair will be conserved in this case). So, the multi-particle interactions should be important in any case Massey and Burhop (1956). On the other hand, one can try to describe them in a simpler approximation, for example, by using the “minimum-image convention”, i.e., summation over the particles located within the distance $\pm L/2$ from the given particle. To the best of our knowledge, the earlier attempts by other authors to employ the simplified summation schemes did not enable them to get the sufficiently stable electron–ion pairs, but this subject still needs to be studied in more detail.
Besides, it should be mentioned that a purely classical electron–ion Coulomb system is not stable in the genuine thermodynamic limit, because the energy is unbounded from below and the trajectories can collapse; so that the quantum-mechanical effects will be essential for the stability of matter. However, this argumentation is irrelevant to the case under consideration: our system cannot reach the thermodynamic equilibrium because it steadily expands up to infinity. In such a case, the most evident stabilizing factor—preventing collapse of the electron trajectories to the singularity—is an approximate conservation (or, more exactly, adiabatic invariance) of the angular momentum $M$ of a captured electron with respect to the corresponding ion, which enters into the effective potential Landau and Lifshitz (1976):
$U_{\rm eff}=\,\frac{M^{2}}{2m_{e}r^{2}}\,-\,\frac{e^{2}}{r}\,.$ (14)
Then, the first term in the right-hand side prevents the electron orbit from the collapse as long as $M$ remains the adiabatic invariant. In other words, the mechanism of (quasi-)stabilization in this case is substantially different from any kinds of quantum corrections, such as a finite electron thermal de Broglie wavelength Demyanov and Levashov (2025), etc.
Let us emphasize that the numerical failures during “head-on” collisions, which were discarded in our simulations, are absolutely irrelevant to the above-mentioned collapse of the classical Coulomb systems in the thermodynamic limit. Really, the discarded runs typically involved the "numerical pumping" of energy (i.e., ejection of the electrons) rather than a collapse of their trajectories to the singularity of the potential. In fact, we have already encountered the same numerical failures in our previous works about the clusterized plasmas Dumin and Lukashenko (2022, 2024). This problem was easily resolved there by using the integration with the adaptive stepsize control (ASSC), particularly, the subroutines from the Numerical Recipes library Press et al. (1992). Unfortunately, these subroutines become unreliable in the case of periodic boundary conditions, which are essential for the present modeling. So, as a simplest remedy, we just discarded the unfavorable initial conditions. In principle, it should be possible to modify the ASSC algorithms for the periodic boundary conditions, but this will require a lot of work on the development of new software.
In fact, a lot of current works employ a much simpler approach, based on the truncated Coulomb interactions; and they successfully reproduce many properties of the ultracold plasmas. However, we were afraid to do so in simulation of the recombination, because the truncated Coulomb potential results in the inherently unclosed classical orbits and, in principle, this can affect the subsequent quantum transitions to the ground atomic states.
@@ V Conclusions
    We suggested a new approach for modeling of the expanding ultracold plasmas, which is based on the “scalable” reference frames, where the effect of expansion formally manifests itself as an effective viscous force in the co-moving coordinates. This enabled us to simulate a non-stationary behavior of such plasmas by a self-consistent way, when all physical parameters (density, temperature, etc.) are determined in the course of time by the same model rather than specified in advance for a set of simulations in the fixed-size boxes.
@@
    To the best of our knowledge, the present work is the first successful ab initio simulation of recombination in the expanding ultracold plasmas, which clearly demonstrates how some of electrons become localized and form the stable compact pairs with the respective ions. Such pairs are assumed to recombine subsequently to the ground atomic states due to the quantum transitions. It is important that we did not employ here any artificial criteria for the formation of classical electron–ion pairs (such as a minimal number of revolutions or a maximal distance between the particles).
@@
    The efficiency of recombination at the realistic experimental conditions obtained by our method turned out to be in agreement both with laboratory measurements and the earlier semi-empirical estimates. For example, if the initial value of the electron coupling parameter ${\Gamma}_{e}$ was about 0.1 or somewhat less, and the plasma cloud expanded by 50–60 times in radius (and, correspondingly, its density dropped by over 5 orders of magnitude), about 20% of the original electrons recombined, which coincides perfectly with the experiment Killian et al. (2001). The earlier numerical modeling Niffenegger, Gilmore, and Robicheaux (2011) resulted in a similar value, 17–18%; but this coincidence should be taken with caution, because the corresponding simulation was carried out in a box of fixed size, and the recombined particles were identified by the semi-empirical criterion.
@@
    Unfortunately, the proposed modeling scheme is rather expensive computationally: for example, a single run takes a few months on the ordinary PC. However, we expect that its efficiency can be improved in future with an optimized software, which should employ the Ewald summation technique for the Coulomb forces and the adaptive stepsize control for integration.
@@ Acknowledgements.
YVD is grateful to A.A. Bobrov, P.R. Levashov, S.A. Mayorov, J.-M. Rost, S.A. Saakyan, U. Saalmann, V.S. Vorob’ev, B.B. Zelener, and B.V. Zelener for fruitful discussions and valuable suggestions as well as to O.E. Stakheev for the assistance in numerical calculations.
@@
Appendix A Examples of the “favorable” and “unfavorable” initial conditions
Figure 7: Temporal variations of the kinetic and potential energies of all particles in the basic cell in the case of “favorable” initial conditions (which are the same as in Figs. 46).
Figure 8: Temporal variations of the kinetic and potential energies in the case of “unfavorable” initial conditions when integration was performed with standard (top panel) and enhanced accuracy (bottom panel).
As was already mentioned in Sec. II.4, the employed integrator was sometimes unable to resolve reliably the close passages of electrons near the ions in the very beginning of simulations, and the respective initial conditions were discarded. Here we provide the corresponding examples.
Figure 7 represents the case of a well-resolved passage of an electron near the pericenter of its strongly-elliptic orbit at the sufficiently large time, when the stable electron–ion pairs were already formed. (The initial conditions are the same as in Figs. 46.) One can see here two major features in variations of the kinetic and potential energies: Firstly, the energetic peaks are symmetric, as should be expected for motion in the elliptic orbit (if weak perturbations from the remote particles are ignored). Secondly, the peaks in kinetic and absolute value of potential energies are approximately of the same magnitude, as should be naturally expected from conservation of the total energy. Really, the potential (Coulomb) energy is actually negative and, in the case of strongly-elliptic orbit, the total energy of the pair is much less than the above-mentioned peak values (i.e., roughly speaking, is close to zero).
An absolutely different situation can be seen in Fig. 8, which corresponds to the case of “unfavorable” initial conditions. In general, a very chaotic pattern of variations is not surprising, because this is an early stage of the plasma cloud expansion, when relaxation did not complete yet and the plasma remained rather hot. One can see in the top panel a sharp asymmetric jump of the kinetic energy at $t^{*}=10.7$, which is evidently unphysical and caused just by numerical errors. Really, this increase in the kinetic energy (about 20 arbitrary units) is an order of magnitude greater than the associated increase in the absolute value of potential energy (about 2 units), which violates the energy conservation condition. Next, the peak of potential energy is approximately symmetric (because position of the electron always changes smoothly with respect to the ion), but the kinetic energy exhibits a strongly asymmetric jump (step). In fact, such cases are well known in UCP simulations and commonly called the “numerical heating”.
However, when the accuracy of numerical integration is substantially improved (bottom panel in Fig. 8), the above-mentioned jump transforms into a tiny symmetric peak, whose amplitude is comparable to the peak of potential energy, as should be expected for a close interparticle collision. Therefore, the problem was purely numerical. Unfortunately, because of the limited computational resources, we were unable to perform the entire set of simulations with the substantially improved accuracy. So, the “unfavorable” cases were just disregarded.